# Holistic Memory Diversification for Incremental Learning in Growing Graphs

### v1

@@ 4 Experiments
This confirms their respective contributions to continual learning. An interesting trend emerges when comparing the individual contributions of the two components. The w/o $\mathcal{L}_{CGSE}$ variant slightly surpasses the performance of w/o $\mathcal{L}_{MISE}$. This suggests that while both components are crucial, adversarial variational embedding learning may have a more pronounced effect in capturing essential and diverse patterns inherent in the data. The best performance occurs with all components, supporting that the proposed components are beneficial individually and collectively, ensuring the model can effectively memorize the previous knowledge while continually adapting to new tasks.
4.1 Related Works
4.1.1 Learning on Growing Graphs.
Graph-based learning often operates under the assumption that the entire graph structure is available upfront. For example, Graph Neural Networks (GNNs) have rapidly become one of the most prominent tools for learning graph-structured data, bridging the gap between deep learning and graph theory. Representative methods includes GCN [29], GraphSAGE [9], and GAT [30], etc. However, these methods predominantly operate on static graphs. Many graphs in real-world applications, such as social networks and transportation systems, are not static but evolve over time. To accommodate this dynamic nature, various methods have been developed to manage growing graph data [31, 32, 33, 34]. For instance, Evolving Graph Convolutional Networks (EvolveGCN) [6] emphasizes temporal adaptability in graph evolution. Temporal Graph Networks (TGNs) [35] operates on continuous-time dynamic graphs represented as a sequence of events. Spatio-Temporal Graph Networks (STGN)[36] integrates spatial and temporal information to enhance prediction accuracy. However, these methods primarily concentrate on a singular task in evolving graphs and often encounter difficulties when more complicated tasks emerge as the graph expands.
4.1.2 Incremental Learning.
Incremental learning [37] refers to a evolving paradigm within machine learning where the model continues to learn and adapt after initial training. The continuous integration of new tasks often leads to the catastrophic forgetting problem. There are usually two types of continual learning settings–class-incremental learning is about expanding the class space within the same task domain, while task-incremental learning involves handling entirely new tasks, which may or may not be related to previous ones [38, 39]. Typically, three different categories of methods have emerged to address the continual learning problem. The first category revolves around regularization techniques [40, 41]. By imposing constraints, these methods prevent significant modifications to model parameters that are critical to previous tasks, ensuring a degree of stability and retention. The second category encompasses parameter-isolation-based approaches [42, 43]. These strategies dynamically allocate new parameters exclusively for upcoming tasks, ensuring that crucial parameters intrinsic to previous tasks remain unscathed. Lastly, memory replay-based methods [44, 45] present a solution by selectively replaying representative data from previous tasks to mitigate the extent of catastrophic forgetting, while are more preferred due to their reduced memory storage requirements and flexibility in parameter training. This paper delve deeper and present an effective strategy for selecting and replaying memory on graphs.
4.1.3 Graph Class-incremental Learning.
@@ 4.1 Related Works
Class-incremental Learning [46] means the specific scenario where the number of classes increases along with the new samples introduced. Graph class-incremental learning [47, 15, 48, 12] specifies this problem to growing graph data–new nodes introduce unseen classes to the graph. Methods of graph class-incremental learning [49, 13, 50, 17, 51] strive to retain knowledge of current classes and adapt to new ones, enabling continuous prediction across all classes. In the past years, various strategies have been proposed to tackle this intricate problem. For example, TWP  [28] employs regularization to ensure the preservation of critical parameters and intricate topological configurations, achieving continuous learning. HPNs [52] adaptively choose different trainable prototypes for incremental tasks. ER-GNN [16] proposes multiple memory sampling strategies designed for the replay of experience nodes. SEM [14] leverages a sparsified subgraph memory selection strategy for memory replay on growing graphs. However, the trade-off between buffer size and replay effect is still a Gordian knot, i.e., aiming for a small buffer size usually results in ineffective knowledge replay. To address this gap, this paper introduces an effective memory selection and replay method that explores and preserve the essential and diversified knowledge contained within restricted nodes, thus improving the model in learning previous knowledge.

### v2

@@ 4 Experiments
The w/o $\mathcal{L}_{CGSE}$ variant slightly surpasses the performance of w/o $\mathcal{L}_{MISE}$. This suggests that while both components are crucial, adversarial variational embedding learning may have a more pronounced effect in capturing essential and diverse patterns inherent in the data. The best performance occurs with all components, supporting that the proposed components are beneficial individually and collectively, ensuring the model can effectively memorize the previous knowledge while continually adapting to new tasks.
5 Related Works
@@
Many graphs in real-world applications, such as social networks and transportation systems, are not static but evolve over time. To accommodate this dynamic nature, various methods have been developed to manage growing graph data (Wang et al., 2020a; Tang & Matteson, 2020; Daruna et al., 2021; Luo et al., 2020). Incremental learning (Parisi et al., 2019; Schwarz et al., 2018; Castro et al., 2018) involves models continuously learning and adapting, often facing the catastrophic forgetting problem. Solutions include regularization techniques (Pomponi et al., 2020; Lin et al., 2023), parameter isolation (Wang et al., 2021a; Lyu et al., 2021), and memory replay (Wang et al., 2021b; Mai et al., 2021). Recent works (Lu et al., 2022; Wang et al., 2022; Yang et al., 2022; Tan et al., 2022) specifies this problem to growing graph data–new nodes introduce unseen classes to the graph. Methods of Graph Incremental Learning (Rakaraddi et al., 2022; Kim et al., 2022; Sun et al., 2023; Su & Wu, 2023; Feng et al., 2023; Liu et al., 2023; Niu et al., 2024) strive to retain knowledge of current classes and adapt to new ones, enabling continuous prediction across all classes. For example, TWP (Liu et al., 2021) employs regularization to ensure the preservation of critical parameters and intricate topological configurations, achieving continuous learning. HPNs (Zhang et al., 2022b) adaptively choose different trainable prototypes for incremental tasks. ER-GNN (Zhou & Cao, 2021) proposes multiple memory sampling strategies designed for the replay of experience nodes. SEM (Zhang et al., 2023) leverages a sparsified subgraph memory selection strategy for memory replay on growing graphs. However, the trade-off between buffer size and replay effect is still a Gordian knot, i.e., aiming for a small buffer size usually results in ineffective memory preserving and knowledge replay. To address this gap, this paper introduces an effective memory selection and replay method that explores and preserves the essential and diversified knowledge contained within restricted nodes, thus improving the model in learning previous knowledge.